# RB-SCD: A New Benchmark for Semantic Change Detection of Roads and Bridges in Traffic Scenes

### v2

@@ V-A Public CD Datasets
In addition to using our proposed dataset, as shown in Tab. III, we have employed three public CD datasets to evaluate the robustness of our model. The LEVIR-CD [7] dataset consists of 637 pairs of RS images with a resolution of 1,024 × 1,024. It mainly focuses on building area changes, such as building growth. The WHU-CD [6] dataset comprises a pair of high-resolution RS images with a size of 32,507 × 15,354 pixels. It covers areas before and after an earthquake and primarily includes building renovations and additions. The SYSU-CD [8] dataset is a collection of 20,000 pairs of high-resolution aerial images, each with a size of 256 × 256 pixels. The images are captured in the Hong Kong area and include urban construction, suburban expansion, and vegetation changes, etc.
TABLE III: Public CD datasets used in the experiments.
Datasets Train/ Val/ Test Image type Space resolution
LEVIR-CD 7120/ 1024/ 2048 Satellite images 0.5 m/pixel
SYSU-CD 12000/ 4000/ 4000 Aerial images 0.5 m/pixel
WHU-CD 5947/ 743/ 744 Aerial images 0.2 m/pixel
@@
TABLE IV: Comparison of quantitative performance for semantic change detection on RB-SCD dataset. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
@@
TABLE V: Performance of the RS foundation model LWGANet on RB-SCD dataset. All metrics are expressed as percentages (%).
@@
TABLE VI: Comparison of quantitative performance for binary change detection on RB-SCD dataset. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
@@
TABLE VII: Comparison of quantitative performance of various CD techniques on public datasets. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance. Results marked with * are obtained from the original papers.
Model LEVIR-CD [7] SYSU-CD [8] WHU-CD [6]
@@
V-B Implementation Details
Our experiments are conducted on a single NVIDIA GeForce RTX 3090 GPU. The AdamW [52] is employed, with a learning rate of 0.0001 and a weight decay parameter of 0.01. Cross entropy loss is used as our loss function, and the maximum number of iterations is set to 40,000. For data augmentation, RandomFlip and PhotoMetricDistortion are applied. The constant $g$ in the ASFF block is set to 0.1, the decay rate $\gamma$ is set to 0.0001, and $\lambda_{0}$ is set to 1. To evaluate the performance of our experiments, we adopt four widely used metrics: recall (Rec), precision (Pre), F1-score (F1) and intersection over union (IoU), where IoU and F1 are the primary metrics, while the others are auxiliary metrics. The following are formulas for defining the evaluation metrics as mentioned above.
$Rec=TP/(TP+FN),$ (13)
$Pre=TP/(TP+FP),$ (14)
$F1=2/(Rec^{-1}+Pre^{-1}),$ (15)
$IoU=TP/(TP+FP+FN),$ (16)
where TP, TN, FP, and FN denote the counts of true positives, true negatives, false positives, and false negatives, respectively.
V-C Quantitative Assessment Results
@@
Results for Semantic Change Detection on RB-SCD Dataset: Tab. IV presents the comparative results of our method and several advanced SCD methods across multiple evaluation metrics. As shown in the table, we have defined 11 types of semantic changes along with one background category, aiming to comprehensively cover the diverse real-world scenarios of road and bridge changes. These change types include both relatively straightforward categories—such as water to bridge and farmland to road, which typically involve simple backgrounds or abundant training samples—and more challenging categories, such as road to vegetation and grass to road, which are characterized by complex backgrounds, ambiguous boundaries, and limited sample sizes. Across this diverse range of semantic changes, our method achieves significant advantages in both the IoU and F1 metrics. Specifically, for the easily detectable categories—water to bridge, bridge to water, water to road, and farmland to road—our model attains IoU scores of 86.83%, 83.58%, 71.68%, and 68.40%, respectively, significantly outperforming other comparison methods. These results highlight the strong capability of our model in semantic discrimination and spatial reasoning. More importantly, our method also performs excellently on several challenging categories, including road to vegetation, road to grass, road to bare land, and grass to road, which suffer from sample scarcity and semantic ambiguity. These types of changes are often affected by similar surface textures, seasonal variability, and complex contextual information. Traditional methods are prone to false positives and missed detections under such conditions, whereas our model consistently demonstrates strong robustness and generalization ability by accurately identifying these subtle semantic transformations.
Furthermore, we evaluate the performance of the latest RS foundation model, LWGANet [47], on our proposed dataset (as shown in Tab. V). The results further underscore the challenging nature of our dataset. By incorporating fine-grained and practically meaningful semantic change categories in traffic scenarios, our dataset provides a more rigorous and instructive benchmark for training and evaluating CD models in real-world applications.
@@
Results for Binary Change Detection on RB-SCD Dataset: As shown in Tab. VI, in the BCD task based on the RB-SCD dataset, different models exhibit varying performance across different evaluation metrics. In terms of the IoU metric, our model ranks first with a score of 65.83%, slightly outperforming ELGC-Net [50], which achieves 65.53%. This indicates that MFDCD demonstrates the best performance in terms of overlap between the predicted and ground truth change regions, enabling more precise localization of change areas. Regarding the F1, our MFDCD model also leads with a score of 79.39%, followed closely by ELGC-Net at 79.18%. Since the F1 is a harmonic mean of Pre and Rec, this result suggests that MFDCD achieves an excellent balance between detection accuracy and completeness. For recall, USSFC-Net [49] achieves the highest score of 75.07% among all models, while our model follows closely with a recall of 74.13%, ranking second. This demonstrates that USSFC-Net has a strong ability to detect all change areas, and our model also exhibits competitive performance in capturing changed regions. In terms of precision, our model outperforms others with a score of 85.47%, while ELGC-Net follows at 85.15%. This implies that our model has a lower false positive rate when identifying change areas. Overall, the results of the BCD task further indicate that existing methods still have room for improvement when applied to our RB-SCD dataset.
@@
Results on Public CD Datasets: To further prove the validity of our model, we also compare our proposed MFDCD with thirteen state-of-the-art CD techniques. To ensure a fair comparison, we reproduce these methods using publicly available code with default parameters. Tab. VII presents the quantitative evaluations of various methods on three public CD datasets, including IoU, F1, Rec, and Pre metrics. As observed in Tab. VII, our method achieves outstanding results across all three publicly available datasets. Specifically, on the WHU-CD dataset, our approach outperforms the second-best method by 2.69% and 1.5% in IoU and F1, respectively. On the SYSU-CD dataset, it achieves improvements of 4.43% and 3.05%, while on the LEVIR-CD dataset, it surpasses the second-best method by 1.06% and 0.62%. These results further demonstrate the effectiveness and robustness of our method.
@@
Figure 7: Qualitative visualization for semantic change detection on RB-SCD dataset. (a) $T_{1}$ images. (b) $T_{2}$ images. (c) HGINet. (d) CdSC. (e) DEFO. (f) ChangeCLIP. (g) Ours. (h) Label.
@@
Figure 9: Heatmaps of our proposed modules. (a) and (g) Input images. (b) and (h) Low-level features from the backbone network. (c) and (i) High-level features from the backbone network. (d) and (j) Feature maps after undergoing the ASFF block. (e) and (k) Feature maps after undergoing the BTFF block. (f) and (l) Feature maps after undergoing the TFF module.
V-D Quantitative Analysis.
As shown in Fig. 7, we compared the visualized results of the proposed MFDCD with other state-of-the-art methods on RB-SCD dataset. From the figure, it can be observed that our method outperforms others in identifying small semantic change regions. For instance, as shown in the third row, our method better distinguishes between changes from grass to road and from vegetation to road—two types of changes with similar characteristics. Regarding pseudo-changes caused by background composition, such as in the fifth row, where a boat passing under a bridge with its bow color similar to the bridge’s edge color, both DEFO [35] and ChangeCLIP [23] made incorrect predictions, while our method accurately detected the changes. Additionally, the proposed dataset presents significant challenges. For example, in the first row, all methods struggle to detect semantic changes under road shadows. As shown in the second row, farmland in autumn and winter appears visually similar to bare land, and in the fourth row, farmland in spring and summer shares similar colors with water, resulting in low inter-class differences. In such challenging scenes, other methods show many false detections, while our method performs slightly better but still leaves room for improvement.
When comparing the BCD results of each method with the ground truth in Fig. 8, it can be observed that for the unchanged areas (black regions representing true negatives), most models show relatively accurate results, with fewer false positives. In contrast, for the changed areas, the performance varies among models. Some models exhibit noticeable missed detections (blue regions representing false negatives), indicating they fail to identify all actual changes. Many models are capable of detecting building—related changes, which are common in the dataset. For example, in areas where buildings are newly constructed or modified, most models can capture at least part of the change, though the precision of boundary detection varies. In regions where there is a transition from bare land to road, our MFDCD and some others show better performance in accurately delineating the change area. Some methods, however, have varying degrees of missed detections. For instance, methods such as BIT [18] and ChangeFormer [19] seem to have less accurate delineation compared to our method and ELGC-Net [50]. In areas where buildings are under construction in one image and completed in another, different models have different performances. Some models, like ICIF [17] and TFI-GR [48], can detect the change but with somewhat blurred boundaries compared to the ground truth (white regions representing true positives). Other models may miss parts of the change. In areas where the change from bare land to road is affected by shadows (resulting in pseudo-changes), most models struggle. Except for our MFDCD, which shows relatively better performance in distinguishing true changes from pseudo-changes caused by shadows, other models either misclassify the shadow areas as changes (false positives in red) or miss the actual changes (false negatives in blue).
TABLE VIII: Ablation studies of DFC and TFF when trained on RB-SCD and public CD datasets. All values are reported as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
@@
TABLE IX: Effectiveness of the DFC module when trained on RB-SCD and public CD datasets. All values are reported in percentage (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
@@
(b) $F_{i}^{1}+H_{i}^{2},i\in\left\{1,2,3,4\right\}$ 55.84 83.52 71.42 89.34
(c) $F_{i}^{t}+LL_{i}^{t},i\in\left\{1,2,3,4\right\}$ 56.10 83.67 71.68 89.45
(d) $F_{i}^{1}+H_{i}^{2},F_{i}^{2}+H_{i}^{1},i\in\left\{1,2\right\}$ 55.70 83.39 71.23 88.90
(e) $F_{i}^{t}+LL_{i}^{t},i\in\left\{3,4\right\}$ 55.86 83.51 71.30 89.24
(f) ASFF 56.42 84.02 72.06 89.98
(g) BTFF 56.87 83.90 71.85 89.77
(h) ASFF+BTFF 57.52 84.69 72.63 90.43
TABLE X: Effectiveness of graph filters with different numbers $k$ in TFF. All values are reported in percentage (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
Method $k$ RB-SCD LEVIR-CD SYSU-CD WHU-CD
@@
TABLE XI: Effectiveness of MFDCD with Different Backbones when Trained on Four CD Datasets. All Values are Reported in Percentage (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
@@
V-E Ablation Study
@@
To demonstrate the effectiveness of our proposed modules, we conduct extensive experiments on RB-SCD and three public datasets. As shown in Tab. VIII, when the DFC module and TFF module are added to the baseline model, the mIoU on RB-SCD dataset increased by 5.44% and 4.78%, respectively. Additionally, the IoU on the LEVIR-CD, SYSU-CD, and WHU-CD datasets improved by 1.6%/2.12%/2.63% and 1.32%/1.91%/1.85%, respectively. As shown in Tab. IX, we conduct detailed ablation studies on the ASFF and BTFF blocks in the DFC module. Initially, we experiment with different coupling methods for ordinary features and frequency domain features. By comparing rows (b) and (c) with row (h), we can observe that ASFF and BTFF effectively dynamically couple visual features with frequency domain features, achieving the best results. Comparing rows (d) and (f), it can be seen that combining the low-level features at time $T_{1}$ with the corresponding high-frequency features at time $T_{2}$, due to semantic mismatches, results in worse performance compared to our designed ASFF. Comparing rows (e) and (g), it is clear that our designed BTFF captures the global structural changes across time, outperforming the coupling of high-level features and low-frequency features at the same time point. As shown in Tab. X, we conduct ablation experiments on the number of layers in the frequency filter of TFF. As the value of $k$ increased, the model’s performance improved. However, when the number of filter layers reached six, the performance stabilized or slightly declined. Subsequent experiments are based on the optimal setting, $k=5$.
@@
To investigate the impact of backbone networks on the performance of the MFDD model in the CD tasks, as shown in Tab.  XI, in terms of parameters and computational complexity, ResNet18 has the lowest cost (13.75M parameters, 11.32G FLOPs), while ResNet50 has the highest (103.78M parameters, 49.02G FLOPs). In terms of detection accuracy, for all four datasets, the deeper the backbone network, the higher the accuracy. However, a deeper network also leads to a doubling of parameters and computational complexity. In this paper, ResNet50 is selected as the backbone network.
@@
To more clearly illustrate the influence of our proposed modules, we visualized heatmaps for these modules. As depicted in columns (b)/(h) and (c)/(i) of Fig. 9, the low-level and high-level features from the backbone network typically exhibit certain characteristics. The low-level features may contain fine-grained but scattered information, while the high-level features, though more abstract, might lose some detailed spatial information. In columns (d)/(j), (e)/(k), and (f)/(l), after undergoing the ASFF block, BTTF block and TFF module respectively, the feature maps show progressive changes. The ASFF block helps to fuse features in a reasonable way, making the information more integrated. The BTTF block further refines the features, enhancing the relevant change information. The TFF module finally outputs feature maps that can better highlight the change regions, eliminating some irrelevant noise and enabling the acquisition of more distinct change-related information, which showcases the efficacy of the introduced modules in feature processing for CD tasks.
VI Conclusion

### v3

@@ 5.1 CD Datasets
In addition to using our proposed dataset, we have employed three public CD datasets to evaluate the robustness of our model. The LEVIR-CD [3] dataset consists of 637 pairs of 1,024 × 1,024 satellite images with a spatial resolution of 0.5 m/pixel. It primarily focuses on building area changes, and is split into 7,120/1,024/2,048 pairs for training, validation, and testing. The SYSU-CD [6] dataset is a collection of 20,000 pairs of 256 × 256 high-resolution aerial images, also at 0.5 m/pixel resolution. Captured in the Hong Kong area, it includes diverse changes like urban construction and suburban expansion, and is divided into 12,000/4,000/4,000 pairs for train/val/test. The WHU-CD [4] dataset comprises a pair of large-format aerial images with a size of 32,507 × 15,354 and a higher resolution of 0.2 m/pixel. It covers areas before and after an earthquake, primarily including building renovations and additions. Following the standard split, it provides 5,947/743/744 image patches.
5.2 Implementation Details
Our experiments are conducted on a single GeForce RTX 3090 GPU. The AdamW is employed, with a learning rate of 0.0001 and a weight decay parameter of 0.01. The maximum number of iterations is set to 40,000. For data augmentation, RandomFlip and PhotoMetricDistortion are applied. The constant $g$ in the ASFF block is set to 0.1, the decay rate $\delta$ is set to 0.0001, and $\lambda_{0}$ is set to 1. We use a weighted cross-entropy loss to mitigate the class imbalance issue in the RB-SCD dataset, which assigns a weight of 1.5 to the losses of under-represented classes, while more common classes retain a weight of 1.0, thereby increasing the penalty for misclassifying rarer change types. To evaluate the performance of our experiments, we adopt four widely used metrics: recall (Rec), precision (Pre), F1-score (F1) and intersection over union (IoU), where IoU and F1 are the primary metrics, while the others are auxiliary metrics.
@@ 5.2 Implementation Details
Table 3: Comparison of quantitative performance for semantic change detection on RB-SCD dataset. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
@@
5.3 Quantitative Assessment Results
Results for Semantic Change Detection on RB-SCD Dataset. Tab. 3 presents the comparative results of our method and several advanced SCD methods across multiple evaluation metrics. Across this diverse range of semantic changes, our method achieves significant advantages in both the IoU and F1 metrics. Specifically, for the easily detectable categories like water->bridge, bridge->water, water-> road and farmland->road, our model attains IoU scores of 86.83%, 83.58%, 71.68%, and
Table 4: Performance of the RS foundation model LWGANet on RB-SCD dataset.
@@ 5.3 Quantitative Assessment Results
Table 5: Comparison of quantitative performance for binary change detection on RB-SCD.
@@
Table 6: Comparison of quantitative performance of various change detection techniques on public datasets. Results marked with * are obtained from the original papers. The ‘-’ symbol indicates that the corresponding data is not available in the source.
Model LEVIR-CD SYSU-CD WHU-CD
@@
68.40%, respectively, significantly outperforming other comparison methods. These results highlight the strong capability of our model in semantic discrimination and spatial reasoning. More importantly, our method also performs excellently on several challenging categories, including road->vegetation, road->grass, road->bare land, and grass->road, which suffer from sample scarcity and semantic ambiguity. These types of changes are often affected by similar surface textures, seasonal variability, and complex contextual information. Traditional methods are prone to false positives and missed detections under such conditions, whereas our model consistently demonstrates strong robustness and generalization ability by accurately identifying these subtle semantic transformations. Furthermore, we evaluate the performance of the latest RS foundation model, LWGANet [38], on our proposed dataset (as shown in Tab. 4). The results further underscore the challenging nature of our dataset.
@@
Results for Binary Change Detection on RB-SCD Dataset. As shown in Tab. 5, in the BCD task based on the RB-SCD dataset, different models exhibit varying performance across different evaluation metrics. In terms of the IoU metric, our model ranks first with a score of 65.83%, slightly outperforming ELGC-Net [19], which achieves 65.53%. This indicates that MFDCD demonstrates the best performance in terms of overlap between the predicted and ground truth change regions, enabling more precise localization of change areas. Regarding the F1, our MFDCD model also leads with a score of 79.39%, followed closely by ELGC-Net at 79.18%. Since the F1 is a harmonic mean of Pre and Rec, this result suggests that MFDCD achieves an excellent balance between detection accuracy and completeness.
@@
Results on Public CD Datasets. To further prove the validity of our model, we also compare our proposed MFDCD with ten state-of-the-art CD models. To ensure a fair comparison, we reproduce these methods using publicly available code with default parameters. As observed in Tab. 6, our method achieves outstanding results across all three publicly available datasets. Specifically, on the WHU-CD dataset, our approach outperforms the second-best method by 2.69% and 1.5% in IoU and F1, respectively. On the SYSU-CD dataset, it achieves improvements of 4.43% and 3.05%, while on the LEVIR-CD dataset, it surpasses the second-best method by 1.06% and 0.62%. These results further demonstrate the effectiveness and robustness of our method.
@@
Figure 9: Qualitative visualization for semantic change detection on RB-SCD. (a) $T_{1}$ images. (b) $T_{2}$ images. (c) HGINet. (d) CdSC. (e) DEFO. (f) ChangeCLIP. (g) Ours. (h) Label.
5.4 Qualitative Analysis.
Qualitative Visualization for BCD Results. Fig. 9 presents a qualitative comparison of MFDCD against state-of-the-art methods on challenging examples from the RB-SCD dataset. The visualizations highlight the tangible benefits of our multimodal, frequency-driven design.
Superior Delineation of Linear Transitions: The first two rows demonstrate the effectiveness of our DFC in modeling linear and heterogeneous change transitions. In the first row, where a new road cuts through grass, other methods produce fragmented or incomplete predictions, especially under the road’s shadow. In contrast, our MFDCD, empowered by the DFC’s ability to model the long-range continuity of linear structures via its BTFF block, generates a much more complete and coherent road segment. Similarly, in the second row (bare land->road), the DFC’s ASFF block, which is sensitive to high-frequency boundary details, allows for a sharper and more precise delineation of the road edges compared to other methods.
Robustness to Semantic Ambiguity: The third and fourth rows showcase the critical role of our TFF in resolving high semantic ambiguity. The third-row case (vegetation->road and grass->road) is a classic example of a scene where multiple land covers transition into a road. While other methods produce a confusing mix of class predictions, our TFF, acting as a semantic reasoning engine, provides a strong contextual prior. This guides the model to a confident and correct interpretation of the scene, resulting in a cleaner and more accurate segmentation. The fourth row further illustrates this, our method correctly identifies the subtle water->road change (light green) that other methods, including the language-guided ChangeCLIP, largely miss, demonstrating the effectiveness of our frequency-domain semantic alignment.
Handling Pseudo-Changes and Complex Scenes: The final row (water->bridge) highlights the robustness of our model. The clean, sharp prediction of the new bridge demonstrates the DFC’s ability to extract the essential structural form while being less sensitive to noise like water texture.
Qualitative Visualization for BCD Results. As shown in Fig. 10, it can be observed that in areas where buildings are newly constructed or modified, most models can capture at least part of the change, though the precision of boundary detection varies. In regions where there is a transition from bare land to road, our MFDCD and some others show better performance in accurately delineating the change area. Some methods, however, have varying degrees of missed detections. For instance, methods such as BIT [16] and ChangeFormer [17] seem to have less accurate delineation compared to our method and ELGC-Net [19]. Some models, like ICIF [14] and TFI-GR [39], can detect the change but with somewhat blurred boundaries compared to the ground truth (white regions representing true positives). Other models may miss parts of the change. In areas where the change from bare land to road is affected by shadows, most models struggle. Except for our MFDCD, which shows relatively better performance in distinguishing true changes from pseudo-changes caused by shadows, other models either misclassify the shadow areas as changes (false positives in red) or miss the actual changes (false negatives in blue).
@@
Table 7: Ablation studies of DFC and TFF when trained on RB-SCD and public CD datasets.
@@
Table 8: Effectiveness of the DFC module when trained on RB-SCD and public CD datasets.
@@
(b) $\mathbf{F}_{i}^{1}+\mathbf{H_{i}^{2}},i\in\left\{1,2,3,4\right\}$ 55.84 83.52 71.42 89.34
(c) $\mathbf{F}_{i}^{1}+\mathbf{H_{i}^{2}},\mathbf{F}_{i}^{2}+\mathbf{H_{i}^{1}},i\in\left\{1,2\right\}$ 55.70 83.39 71.23 88.90
(d) $\mathbf{F}_{i}^{t}+\mathbf{LL}_{i}^{t},i\in\left\{3,4\right\}$ 55.86 83.51 71.30 89.24
(e) ASFF 56.42 84.02 72.06 89.98
(f) BTFF 56.87 83.90 71.85 89.77
(g) TFIM +TFF 55.98 83.86 71.63 88.97
(h) ASFF+BTFF (DFC) 57.52 84.69 72.63 90.43
Table 9: Effectiveness of the TFF module when trained on RB-SCD and public CD datasets.
Method RB-SCD LEVIR-CD SYSU-CD WHU-CD
(a) - 55.89 83.25 71.10 89.28
(b) DFT+MLP 56.31 83.59 71.42 89.57
(c) DFT+1D CNN 56.47 83.73 71.62 89.68
(d) DFT+GAT 56.67 83.93 71.69 89.75
(e) CLIP Embedding + Concat 56.38 83.54 71.85 89.84
(f) DFT + Graph Filter (TFF) 57.52 84.69 72.63 90.43
Table 10: Effectiveness of graph filters with different numbers $l$ in TFF.
Method $l$ RB-SCD LEVIR-CD SYSU-CD WHU-CD
@@
Table 11: Effectiveness of MFDCD with different backbones when trained on RB-SCD and public CD datasets.
@@
5.5 Ablation Study
To demonstrate the effectiveness of our proposed modules, we conduct extensive experiments on RB-SCD and three public datasets.
Effectiveness of DFC and TFF Modules. As shown in Tab. 7, adding the DFC and TFF to the baseline model increases the mIoU on the RB-SCD dataset by 5.44% and 4.78%, respectively. Additionally, the IoU on the LEVIR-CD, SYSU-CD, and WHU-CD datasets improved by 1.6%/2.12%/2.63% and 1.32%/1.91%/1.85%, respectively.
@@ 5.5 Ablation Study
Detailed Analysis on DFC Module. As shown in Tab. 8, we conduct detailed ablation studies on the ASFF and BTFF blocks in the DFC module. Initially, we experiment with different coupling methods for ordinary features and frequency domain features. By comparing rows (b) and (c) with row (h), we can observe that ASFF and BTFF effectively dynamically couple visual features with frequency domain features, achieving the best results. Comparing rows (c) and (e), it can be seen that combining the low-level features at time $T_{1}$ with the corresponding high-frequency features at time $T_{2}$, due to semantic mismatches, results in worse performance compared to our designed ASFF. Comparing rows (d) and (f), it is clear that our designed BTFF captures the global structural changes across time, outperforming the coupling of high-level features and low-frequency features at the same time point. To further validate our design, we replace our DFC with TFIM [39] (row (g)), a strong module for spatial feature interaction. Although TFIM provides a notable improvement over the baseline, our frequency-domain DFC yields better performance, demonstrating the superiority of explicitly modeling structural features in the frequency domain for this task.
Detailed Analysis on TFF Module. As shown in Tab. 9, the MLP (row (b)), which processes each frequency component independently, performs the worst, highlighting the necessity of modeling relational information. The 1D CNN (row (c) captures only local relationships and offers a slight improvement. Notably, while the GAT adaptively learns edge weights (row (d)), our proposed method (row (e)) with a simple graph filter over a fixed fully connected graph still achieves the best results. This indicates long-range dependency across the entire spectrum is more effective than attempting to dynamically prune or re-weight connections To validate our frequency-domain approach, we benchmark our TFF module against a powerful CLIP Embedding + Concat [14] mechanism (row (e)) that fuses semantics in the spatial/token domain. It shows that while this mechanism improves upon the vision-only baseline, our TFF-based model (row (f)) achieves superior performance. This suggests that extracting a robust, global semantic signature via frequency analysis is more effective for disambiguating complex RS scenes than performing direct spatial alignment with text tokens. As shown in
Figure 11: Heatmaps of our proposed modules. (a) and (g) Input images. (b) and (h) Low-level features from the backbone. (c) and (i) High-level features from the backbone network. (d) and (j) Feature maps after undergoing the ASFF block. (e) and (k) Feature maps after undergoing the BTFF block. (f) and (l) Feature maps after undergoing the TFF module.
Figure 12: Hyperparameter sensitivity analysis of the dynamic sparsity constraint in ASFF. We evaluate the impact of the initial sparsity $\lambda_{0}$ (left), the decay rate $\delta$ (middle) and the constant $g$ (right) on the mIoU performance on RB-SCD.
Tab. 10, we conduct ablation experiments on the number of layers in the frequency filter of TFF. As the value of $l$ increased, the model’s performance improved. However, when the number of filter layers reached six, the performance stabilized or slightly declined. Subsequent experiments are based on the optimal setting, $l=5$.
@@
Backbone Analysis. To investigate the impact of backbone networks of the MFDCD model, as shown in Tab. 11, in terms of parameters and computational complexity, ResNet18 has the lowest cost, while ResNet50 has the highest. In terms of detection accuracy, for all four datasets, the deeper the backbone network, the higher the accuracy. However, a deeper network also leads to a doubling of parameters and computational complexity. In this paper, ResNet50 is selected as the backbone network.
@@
Visualization Analysis. To more clearly illustrate the influence of our proposed modules, we visualized heatmaps for these modules. As depicted in columns (b)/(h) and (c)/(i) of Fig. 11, the low-level and high-level features from the backbone network typically exhibit certain characteristics. The low-level features may contain fine-grained but scattered information, while the high-level features, though more abstract, might lose some detailed spatial information. In columns (d)/(j), (e)/(k), and (f)/(l), after undergoing the ASFF block, BTTF block and TFF module respectively, the feature maps show progressive changes. The ASFF block helps to fuse features in a reasonable way, precisely delineate the boundaries. The BTTF block further maintains the continuity and integrity of linear structures. The TFF module finally outputs feature maps that can better highlight the change regions, eliminating some irrelevant noise and resolving semantic ambiguities.
Hyperparameter Analysis. As shown in Fig. 12, we conduct a sensitivity analysis on the key hyperparameters $\lambda_{0}$, $\delta$ and $g$ of our dynamic sparsity constraint. The performance remains stable across a reasonable range of values, demonstrating the robustness of our approach. Our chosen values ($\lambda_{0}$=1.0, $\delta$=0.0001,gg=0.1) provide the best performance. This dynamic mechanism proves more effective than a fixed threshold, as it adapts the level of detail refinement to the model’s learning stage, ensuring both stability in early training and precision in later stages.
6 Conclusion